### v1

# Coherence length in a dilute flat-band superconductor

Abstract
To explore the influence of quantum-geometric effects on the Ginzburg-Landau coherence length in a dilute flat-band superconductor, we adopt a BCS-BEC crossover approach to the multiband pyrochlore-Hubbard model near the critical temperature for superconductivity. Our self-consistent formulation for this three-dimensional lattice benchmarks very well against the so-called zero-temperature coherence length, demonstrating the monotonic decay of the coherence length to zero as the interaction strength increases. Additionally, we show that the effective mass of the many-body bound states (i.e., Cooper pairs) is nearly identical to that of the lowest-lying two-body bound states in the dilute flat-band limit.
I Introduction
Recent theoretical studies have uncovered a surprising link between the quantum geometry of Bloch states and superconductivity in multiband systems, significantly advancing our understanding of the Cooper-pairing mechanism in flat-band systems [1, 2]. Under certain restrictive conditions, such as time-reversal symmetry and uniform pairing, it has been revealed that the quantum-metric tensor plays a critical role in determining key observables. These include the superfluid weight, superfluid density, critical transition temperature, low-energy collective Goldstone and Leggett modes, Ginzburg-Landau (GL) coherence length, London penetration depth, GL parameter, and upper critical magnetic field  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The common factor among these observables is the effective mass of the superfluid carriers [12, 13]. Despite the infinite band mass of particles in the underlying flat band, even an infinitesimal interaction gives Cooper pairs a finite effective mass through virtual interband transitions mediated by the quantum metric. This mechanism facilitates the emergence of superconductivity in a flat band in the presence of other bands [14]. Furthermore, the geometric origin is most evident in the effective mass of the lowest-lying two-body bound states, as demonstrated by exact calculations [15].
In multiband superconductors, geometric effects, while inherent, are generally overshadowed by conventional effects in the BCS limit [1, 2]. However, as the bandwidth of a Bloch band narrows, conventional effects diminish due to their dependence on the derivative of the Bloch bands. In contrast, geometric effects, which are determined by the derivative of the Bloch states, can become dominant under these conditions. Inspired by recent technological advances in creating two-dimensional materials with nearly-flat electronic bands [16], there is a growing surge of interest in understanding geometric effects in multiband superconductors [17, 18, 19]. Among them, there has been a perplexing GL study regarding the size of Cooper pairs in a flat-band superconductor, suggesting they are limited by a fundamental length scale determined by the quantum geometry of the flat band [20]. However, subsequent work using the Bogoliubov-de Gennes approach for various lattice models in one and two dimensions has demonstrated that characteristic correlation lengths can be smaller than one lattice spacing without being constrained by the quantum metric [21].
To address these discrepancies [20, 21], we investigate the quantum-geometric effects on the GL coherence length in a dilute flat-band superconductor using a BCS-BEC crossover approach to the multiband pyrochlore-Hubbard model near the critical temperature for pairing. Our self-consistent results are in good agreement with the zero-temperature coherence length, demonstrating that GL coherence length decreases to zero monotonously as the interaction strength increases. In addition, we find that the effective mass of Cooper pairs aligns closely with that of the lowest-lying two-body bound states in the dilute regime. The pyrochlore lattice, akin to well-studied two-dimensional models like the Mielke-checkerboard and kagome lattices known for uniform pairing, provides an ideal framework for exploring three-dimensional flat-band superconductivity [6]. Recent demonstrations of flat bands and superconductivity in materials like pyrochlore metal CaNi2 [18] and pyrochlore superconductor CeRu2 [19] underscore the relevance of this model.
The rest of of the paper is organized as follows. In Sec. II, we introduce the multiband Hubbard Hamiltonian in reciprocal space, review the effective-action formalism near $T_{c}$, and derive the expansion coefficients for the time-dependent-GL (TDGL) theory. In Sec. III, we discuss self-consistency relations that are used to study dilute flat-band superconductivity in the pyrochlore lattice. In Sec. IV, we introduce the zero-temperature coherence length. In Sec. V, we present the numerical results and analyze them in various limits. The paper ends with a summary and outlook in Sec. VI, and the breakdown of zero-temperature coherence length is discussed in the Appendix.
II TDGL Theory
Having the pyrochlore-Hubbard model that we recently studied at zero temperature ($T=0$) in mind [6], we begin by introducing the multiband Hubbard model with an onsite attractive interaction and generic hopping terms. We then review the derivation of the effective action near the critical temperature $T_{c}$ for superconductivity and apply the resultant TDGL theory to the pyrochlore lattice, which exhibits time-reversal symmetry and uniform pairing.
II.1 Multiband Hubbard Hamiltonian
Just as the kagome lattice is a line graph of the honeycomb lattice and the Mielke checkerboard lattice is a line graph of the square lattice, both of which feature a flat tight-binding band in two dimensions, the pyrochlore lattice is a line graph of the diamond lattice and it features two degenerate flat bands in three dimensions. Its crystal structure consists of a face-centered-cubic Bravais lattice with a four-point basis, leading to a truncated-octahedron shaped Brillouin Zone (BZ) [6]. This is in such a way that the total number of lattice sites is $N=4N_{c}$, where $N_{c}=\sum_{\mathbf{k}\in\mathrm{BZ}}1$ counts the number of unit cells in real space, with $\mathbf{k}=(k_{x},k_{y},k_{z})$ denoting the crystal momentum in units of $\hbar\to 1$.
Upon Fourier transformation from the site representation to the reciprocal space, the multiband Hubbard model becomes
$\mathcal{H}$ $=\sum_{SS^{\prime}\mathbf{k}\sigma}\psi_{S\mathbf{k}\sigma}^{\dagger}(h_{SS^{\prime}}^{\mathbf{k}\sigma}-\mu_{\sigma}\delta_{SS^{\prime}})\psi_{S^{\prime}\mathbf{k}\sigma}$
$-\frac{U}{N_{c}}\sum_{S\mathbf{k}\mathbf{k^{\prime}}\mathbf{q}}\psi_{S\mathbf{k}\uparrow}^{\dagger}\psi_{S,-\mathbf{k}+\mathbf{q},\downarrow}^{\dagger}\psi_{S,-\mathbf{k^{\prime}}+\mathbf{q},\downarrow}\psi_{S\mathbf{k^{\prime}}\uparrow},$ (1)
where $\psi_{S\mathbf{k}\sigma}^{\dagger}$ creates a spin $\sigma=\{\uparrow,\downarrow\}$ particle on the sublattice $S=\{A,B,C,D\}$ and $\delta_{ij}$ is a Kronecker delta. Here, $U\geq 0$ is the strength of the attractive onsite interaction between $\uparrow$ and $\downarrow$ particles, and $\mu_{\sigma}$ is the chemical potential determining their average numbers in the ensemble. In the specific case of the pyrochlore lattice, which is our primary interest in this paper, the matrix elements of the Bloch Hamiltonian $\mathbf{h}_{\mathbf{k}\sigma}$ can be written as $h_{SS}^{\mathbf{k}\sigma}=0,$ $h_{AB}^{\mathbf{k}\sigma}=-2\bar{t}\cos\big{(}\frac{k_{y}+k_{z}}{4}a\big{)},$ $h_{AC}^{\mathbf{k}\sigma}=-2\bar{t}\cos\big{(}\frac{k_{x}+k_{z}}{4}a\big{)},$ $h_{AD}^{\mathbf{k}\sigma}=-2\bar{t}\cos\big{(}\frac{k_{x}+k_{y}}{4}a\big{)},$ $h_{BC}^{\mathbf{k}\sigma}=-2\bar{t}\cos\big{(}\frac{k_{x}-k_{y}}{4}a\big{)},$ $h_{BD}^{\mathbf{k}\sigma}=-2\bar{t}\cos\big{(}\frac{k_{x}-k_{z}}{4}a\big{)}$ and $h_{CD}^{\mathbf{k}\sigma}=-2\bar{t}\cos\big{(}\frac{k_{y}-k_{z}}{4}a\big{)},$ where $\bar{t}$ is the tight-binding hopping parameter between the nearest-neighbor sites and $a$ is the side-length of the conventional simple-cubic cell [6]. Thus, the Bloch bands are determined by $\sum_{S^{\prime}}h_{SS^{\prime}}^{\mathbf{k}\sigma}n_{S^{\prime}\mathbf{k}\sigma}=\varepsilon_{n\mathbf{k}\sigma}n_{S\mathbf{k}\sigma},$ leading to $\varepsilon_{1\mathbf{k}\sigma}=-2\bar{t}(1+\sqrt{1+\alpha_{\mathbf{k}}})$ and $\varepsilon_{2\mathbf{k}\sigma}=-2\bar{t}(1-\sqrt{1+\alpha_{\mathbf{k}}})$ for the dispersive bands, where $\alpha_{\mathbf{k}}=\cos(k_{x}a/2)\cos(k_{y}a/2)+\cos(k_{y}a/2)\cos(k_{z}a/2)+\cos(k_{x}a/2)\cos(k_{z}a/2),$ and $\varepsilon_{3\mathbf{k}\sigma}=\varepsilon_{4\mathbf{k}\sigma}=2\bar{t}$ for the degenerate flat bands. Note that $\varepsilon_{2\mathbf{k}\sigma}$ touches the flat bands at $\mathbf{k=0}$. To ensure that these flat bands appear at the bottom of the Bloch spectrum, we set $\bar{t}\to-t$ and choose $t>0$ as the unit of energy. This allows us to construct a relatively simple BCS-BEC crossover formalism for a dilute flat-band superconductor near $T_{c}$, and study some of its properties as functions of $U$. Here, $n_{S\mathbf{k}\sigma}=\langle S|n_{\mathbf{k}\sigma}\rangle$ is the sublattice projection of the Bloch state $|n_{\mathbf{k}\sigma}\rangle$, which plays an important role throughout this paper.
II.2 Effective action near $T\leq T_{c}$
In order to construct a microscopic TDGL theory near $T_{c}$ [22], we first extract the effective free-energy density of the system per lattice site as a function of the pairing order parameter from the effective action, i.e., from $\frac{T}{N}(S_{2}+S_{4}),$ where $T$ is the temperature in units of the Boltzmann constant $k_{\mathrm{B}}\to 1$, and $S_{2}$ ($S_{4}$) is the quadratic (quartic) contribution to the action [5]. The saddle-point action $S_{0}$ is not needed here because it does not depend on the order parameter at $T_{c}$, and $S_{1}$ and $S_{3}$ vanish at any $T$ due to the saddle-point condition.
Using the Grassmann functional-integral formalism, one can show that $\mathcal{S}_{2}=\frac{N_{c}}{T}\sum_{SS^{\prime}q}\Lambda_{S}^{*}(q)\Gamma^{-1}_{SS^{\prime}}(q)\Lambda_{S^{\prime}}(q)$  [5], where $q\equiv(\mathbf{q},\mathrm{i}\nu_{\ell})$ is a collective index with $\nu_{\ell}=2\ell\pi T$ denoting the bosonic Matsubara frequency, $\Lambda_{S}(q)$ plays the role of the pairing order parameter for sublattice $S$, and
$\Gamma_{SS^{\prime}}^{-1}(q)=\frac{\delta_{SS^{\prime}}}{U}$ $+\frac{1}{2N_{c}}\sum_{nm\mathbf{k}}\frac{\mathcal{X}_{n\mathbf{k}\uparrow}+\mathcal{X}_{m,-\mathbf{k}+\mathbf{q},\downarrow}}{\mathrm{i}\nu_{\ell}-\xi_{n\mathbf{k}\uparrow}-\xi_{m,-\mathbf{k}+\mathbf{q},\downarrow}}$
×$\times n_{S\mathbf{k}\uparrow}n_{S^{\prime}\mathbf{k}\uparrow}^{*}m_{S^{\prime},-\mathbf{k}+\mathbf{q},\downarrow}^{*}m_{S,-\mathbf{k}+\mathbf{q},\downarrow},$ (2)
is the matrix element of the inverse pair-fluctuation propagator $\boldsymbol{\Gamma}^{-1}(q).$ Here, $\mathcal{X}_{n\mathbf{k}\sigma}=\tanh\big{(}\frac{\xi_{n\mathbf{k}\sigma}}{2T}\big{)}$ is a thermal factor and $\xi_{n\mathbf{k}\sigma}=\varepsilon_{n\mathbf{k}\sigma}-\mu_{\sigma}.$ This propagator suggests that the generalized Thouless condition $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},0)=0$ determines $T_{c}$ of a multiband Hubbard model for any center-of-mass momentum $\mathbf{q}$. For instance, $T_{c}=\max\{T_{c_{1}},T_{c_{2}},T_{c_{3}},T_{c_{4}}\}$ in the case of the pyrochlore lattice, where $T_{c_{j}}$ is determined by setting the $j$th eigenvalue of $\boldsymbol{\Gamma}^{-1}(\mathbf{q},0)$ to 00. In our numerical calculations, we observed that $T_{c_{1}}>T_{c_{2}}=T_{c_{3}}=T_{c_{4}}$ for any given set of $(\mu_{\uparrow}=\mu_{\downarrow},U)$ in the $\mathbf{q}\to\mathbf{0}$ limit. Furthermore, we observed that the associated eigenvector of $\boldsymbol{\Gamma}^{-1}(\mathbf{q},0)$ that corresponds to the $T_{c_{1}}$ solution is uniform in a unit cell, i.e., $[\Lambda_{A}(\mathbf{q}),\Lambda_{B}(\mathbf{q}),\Lambda_{C}(\mathbf{q}),\Lambda_{D}(\mathbf{q})]=\Lambda_{0}(\mathbf{q})[1,1,1,1]$ in the $\mathbf{q}\to\mathbf{0}$ limit. It is pleasing to note that the latter observation aligns perfectly with our previous finding that the uniform-pairing condition is satisfied exactly for the lowest-lying two-body bound states of the pyrochlore lattice when $\mathbf{q\to 0}$ [6].
Motivated by these numerical insights and to facilitate further analytical progress, we next adopt the following assumptions that are satisfied by the pyrochlore lattice. ($i$) The lattice manifests time-reversal symmetry, leading to $n_{S,-\mathbf{k},\downarrow}^{*}=n_{S\mathbf{k}\uparrow}=n_{S\mathbf{k}}$ and $\xi_{n,-\mathbf{k},\downarrow}=\xi_{n\mathbf{k}\uparrow}=\xi_{n\mathbf{k}}.$ ($ii$) The low-$q$ order parameters are uniform in a unit cell, leading to $\Lambda_{0}(q)=\Lambda_{S}(q)$ for all sublattices. Under these assumptions, the quadratic action can be written as $\mathcal{S}_{2}=\frac{N}{T}\sum_{q}\Gamma^{-1}_{0}(q)|\Lambda_{0}(q)|^{2}$  [5], where
$\Gamma^{-1}_{0}(q)$ $=\frac{1}{U}+\frac{1}{2N}\sum_{nm\mathbf{k}}\frac{\mathcal{X}_{n\mathbf{k}}+\mathcal{X}_{m,\mathbf{k-q}}}{\mathrm{i}\nu_{\ell}-\xi_{n\mathbf{k}}-\xi_{m,\mathbf{k-q}}}|\langle n_{\mathbf{k}}|m_{\mathbf{k-q}}\rangle|^{2}.$ (3)
Thus, the Thouless condition reduces to $\Gamma_{0}^{-1}(\mathbf{0},0)=0$ for stationary BCS-type pairing, determining $T_{c}$ of a uniformly-paired multiband Hubbard model in the presence of time-reversal symmetry. Similarly, by making use of the Grassmann functional-integral formalism, and under the same assumptions discussed above, one can approximate that $\mathcal{S}_{4}=\frac{Nb_{0}}{2T}\sum_{q_{1}q_{2}q_{3}}\Lambda_{0}^{*}(q_{1}+q_{2}+q_{3})\Lambda_{0}(q_{1})\Lambda_{0}^{*}(-q_{2})\Lambda_{0}(q_{3})$  [5]. Here, a positive $b_{0}$ coefficient not only guarantees the energetic stability of the TDGL theory, but it also characterizes the repulsive interaction between Cooper pairs as discussed next.
II.3 Coefficients for the TDGL expansion
When $T\lesssim T_{c}$, the coefficients for the microscopic TDGL theory can be determined through the expansion of the inverse propagator $\Gamma^{-1}_{0}(\mathbf{q},\mathrm{i}\nu_{\ell}\to\omega+\mathrm{i}0^{+})=-a_{0}\epsilon(T)+\frac{1}{2}\sum_{ij}c_{ij}q_{i}q_{j}+d\omega+\cdots$ in the low-$(\mathbf{q},\omega)$ regime, where $\epsilon(T)=(T_{c}-T)/T_{c}$ changes sign across $T_{c}$ in accordance with the Landau theory of second-order phase transitions [5, 22]. This expansion leads to
$a_{0}$ $=\frac{1}{N}\sum_{n\mathbf{k}}\bigg{[}\frac{\mathcal{Y}_{n\mathbf{k}}}{4T_{c}}+\frac{\partial\mu}{\partial T}\bigg{(}\frac{\mathcal{Y}_{n\mathbf{k}}}{4\xi_{n\mathbf{k}}}-\frac{T_{c}\mathcal{X}_{n\mathbf{k}}}{2\xi_{n\mathbf{k}}^{2}}\bigg{)}\bigg{]},$ (4)
$b_{0}$ $=\frac{1}{N}\sum_{n\mathbf{k}}\left(\frac{\mathcal{X}_{n\mathbf{k}}}{4\xi_{n\mathbf{k}}^{3}}-\frac{\mathcal{Y}_{n\mathbf{k}}}{8T_{c}\xi_{n\mathbf{k}}^{2}}\right),$ (5)
$c_{ij}^{\mathrm{intra}}$ $=\frac{1}{N}\sum_{n\mathbf{k}}\left(\frac{\mathcal{X}_{n\mathbf{k}}}{4\xi_{n\mathbf{k}}^{3}}-\frac{\mathcal{Y}_{n\mathbf{k}}}{8T_{c}\xi_{n\mathbf{k}}^{2}}\right)\dot{\xi}_{n\mathbf{k}}^{i}\dot{\xi}_{n\mathbf{k}}^{j},$ (6)
$c_{ij}^{\mathrm{inter}}$ $=\frac{1}{N}\sum_{n\mathbf{k}}\frac{\mathcal{X}_{n\mathbf{k}}}{2\xi_{n\mathbf{k}}}g^{n\mathbf{k}}_{ij}-\frac{1}{2N}\sum_{n,m\neq n,\mathbf{k}}\frac{\mathcal{X}_{n\mathbf{k}}+\mathcal{X}_{m\mathbf{k}}}{\xi_{n\mathbf{k}}+\xi_{m\mathbf{k}}}g^{nm\mathbf{k}}_{ij},$ (7)
$d$ $=\frac{1}{N}\sum_{n\mathbf{k}}\frac{\mathcal{X}_{n\mathbf{k}}}{4\xi_{n\mathbf{k}}^{2}}+\frac{\mathrm{i}\pi}{8T_{c}N}\sum_{n}D_{n}(\mu)\theta_{\mu},$ (8)
where $\mathcal{Y}_{n\mathbf{k}}=\mathrm{sech^{2}}\big{(}\frac{\xi_{n\mathbf{k}}}{2T}\big{)}$ is a thermal factor and $\dot{\xi}_{n\mathbf{k}}^{i}=\partial\xi_{n\mathbf{k}}/\partial k_{i}.$ In order to reproduce the correct $U_{pp}$ that is previously obtained from the low-energy collective-mode analysis at $T=0$ [6], i.e., see Sec. V for further discussion, here the expansion coefficients are given per lattice site. We note that all of these coefficients must be evaluated at $T_{c}$ self-consistently with $\mu$. In addition, since the factor $\frac{\partial\mu}{\partial T}$ plays a crucial role away from the BCS limit, Eq. (4) has to be handled with care in flat-band superconductors, which is described in Sec. V. However, numerical implementation of the rest of the coefficients is a straightforward task once $(\mu,T_{c})$ is computed for a desired $U$.
Motivated by the recent surge of interest in the quantum-geometric effects, here we also split the kinetic coefficient into two contributions $c_{ij}=c_{ij}^{\mathrm{intra}}+c_{ij}^{\mathrm{inter}},$ depending on whether the intraband or interband processes are involved. Note that, in comparison to our previous work [5], here we used integration by parts $\sum_{\mathbf{k}}\mathcal{X}_{n\mathbf{k}}\ddot{\xi}_{n\mathbf{k}}^{ij}/\xi_{n\mathbf{k}}^{2}=\sum_{\mathbf{k}}\big{[}2\mathcal{X}_{n\mathbf{k}}/\xi_{n\mathbf{k}}^{3}-\mathcal{Y}_{n\mathbf{k}}/(2T\xi_{n\mathbf{k}}^{2})\big{]}\dot{\xi}_{n\mathbf{k}}^{i}\dot{\xi}_{n\mathbf{k}}^{j}$ and $\sum_{\mathbf{k}}\mathcal{Y}_{n\mathbf{k}}\ddot{\xi}_{n\mathbf{k}}^{ij}/\xi_{n\mathbf{k}}=\sum_{\mathbf{k}}\big{[}\mathcal{Y}_{n\mathbf{k}}/\xi_{n\mathbf{k}}^{2}+\mathcal{X}_{n\mathbf{k}}\mathcal{Y}_{n\mathbf{k}}/(T\xi_{n\mathbf{k}})\big{]}\dot{\xi}_{n\mathbf{k}}^{i}\dot{\xi}_{n\mathbf{k}}^{j},$ and reexpressed $c_{ij}^{\mathrm{intra}}$ in the new form given by Eq. (6), where $\ddot{\xi}_{n\mathbf{k}}^{ij}=\partial^{2}\xi_{n\mathbf{k}}/\partial k_{i}\partial k_{j}.$ This coefficient is the so-called conventional contribution, as it is simply a sum over its single-band counterpart [5]. Similarly, the coefficients $a_{0}$, $b_{0}$ and $d$ are sums over their single-band counterparts. In contrast, the geometric contribution $c_{ij}^{\mathrm{inter}}$ is controlled by the quantum-metric tensor $g_{ij}^{n\mathbf{k}}=\sum_{n\neq m}g_{ij}^{nm\mathbf{k}}$ of the $n$th Bloch band [23, 24, 25] and its band-resolved quantum-metric tensor $g_{ij}^{nm\mathbf{k}}=2\mathrm{Re}\left[\langle\dot{n}_{\mathbf{k}}^{i}|m_{\mathbf{k}}\rangle\langle m_{\mathbf{k}}|\dot{n}_{\mathbf{k}}^{j}\rangle\right],$ where $\mathrm{Re}$ denotes the real part and $|\dot{n}_{\mathbf{k}}^{i}\rangle=\partial|n_{\mathbf{k}}\rangle/\partial k_{i}.$ Such terms follow from the low-$\mathbf{q}$ expansion of the Bloch factor
$|\langle n_{\mathbf{k}}|m_{\mathbf{k-q}}\rangle|^{2}=\delta_{nm}-\frac{1}{2}\sum_{ij}[g_{ij}^{n\mathbf{k}}\delta_{nm}+g_{ij}^{nm\mathbf{k}}(\delta_{nm}-1)]q_{i}q_{j}$ (9)
that appears in Eq. (3). It is important to highlight that $c_{ij}^{\mathrm{inter}}$ does not have any contribution from the band touchings, i.e., the first term of Eq. (7) cancels those touching contributions from the second term whenever $\xi_{n\mathbf{k}}=\xi_{m\mathbf{k}}$ for any $n\neq m$. For instance, in the case of pyrochlore lattice, there is no inter-flat-band contribution to $c_{ij}^{\mathrm{inter}}$ among the degenerate flat bands as they touch each other everywhere in the BZ.
Furthermore, we note that the dynamic coefficient $d$ is a complex number in general [5, 22], and its imaginary part is determined by the density of states $D_{n}(\varepsilon)=\sum_{\mathbf{k}}\delta(\varepsilon-\varepsilon_{n\mathbf{k}})$ of the $n$th Bloch band where $\theta_{\varepsilon}=\theta(\varepsilon-\min\{\varepsilon_{n\mathbf{k}}\})\theta(\max\{\varepsilon_{n\mathbf{k}}\}-\varepsilon).$ Here, $\delta(x)$ is the Dirac-delta distribution and $\theta(x)$ is the Heaviside-step function. Thus, $d$ has a dominant positive imaginary part in the BCS limit when $\mu$ lies within any one of the Bloch bands. This is nothing but a reflection of the continuum of fermionic excitations into which a Cooper pair can decay, suggesting that the dynamics of the order parameter is overdamped in the BCS limit. In sharp contrast, $d$ becomes purely real in the BEC limit suggesting that the dynamics of the order parameter is propagating. For the dilute flat-band superconductor of interest in this paper, the latter turns out to the case for any $U\neq 0$.
Assuming this is the case, and by making an analogy with the Gross-Pitaevskii equation for a weakly-interacting atomic Bose gas, i.e., through a scaling of the order parameter [5, 22], we define $\mu_{p}(T)=a_{0}\epsilon(T)/d$ as the effective chemical potential of the Cooper pairs, $(M_{p}^{-1})_{ij}=c_{ij}/d$ as the inverse effective-mass tensor $\mathbf{M}_{p}^{-1}$ of the pairs, and $U_{pp}=b_{0}/d^{2}$ as the effective onsite repulsive interaction between the pairs. This identification suggests that $\mathbf{M}_{p}$ is composed of a conventional contribution $(M_{p}^{-1})^{\mathrm{intra}}_{ij}=c_{ij}^{\mathrm{intra}}/d$ and a geometric one $(M_{p}^{-1})^{\mathrm{inter}}_{ij}=c_{ij}^{\mathrm{inter}}/d.$ In this paper, we are primarily interested in the GL coherence length $(\xi_{T}^{2})_{ij}=(\xi_{\mathrm{GL}}^{2})_{ij}/\epsilon(T),$ whose temperature-independent prefactor is given by
$(\xi_{\mathrm{GL}}^{2})_{ij}$ $=\frac{c_{ij}}{2a_{0}}.$ (10)
This relation again suggests that $\xi_{\mathrm{GL}}$ is composed of a conventional contribution $(\xi_{\mathrm{GL}}^{2})^{\mathrm{intra}}_{ij}=c_{ij}^{\mathrm{intra}}/(2a_{0})$ and a geometric one $(\xi_{\mathrm{GL}}^{2})^{\mathrm{inter}}_{ij}=c_{ij}^{\mathrm{inter}}/(2a_{0}).$ Since the calculation of $\xi_{\mathrm{GL}}$ requires $\mu$ and $T_{c}$ as inputs, we next describe our recipe for their self-consistent evaluation as a function of $U$, which is appropriate for a dilute flat-band superconductor in the pyrochlore lattice.
III Self-consistency relations: $(\mu,T_{c})$
As the simplest route, we follow the usual finite-temperature BCS-BEC crossover formalism [22, 26] to construct a self-consistent theory based on the Thouless condition, which is equivalent to both the saddle-point condition and the mean-field order-parameter equation, and the number equation $\mathcal{N}=-\partial\Omega/\partial\mu.$ Here, $\Omega$ is the thermodynamic potential, and keeping its corrections at the Gaussian order, i.e., $\Omega_{\mathrm{G}}=\Omega_{0}+\Omega_{2},$ is known to be sufficient in producing a qualitatively correct physical description of the system for all $U\neq 0$.
One can show quite generally that $\Omega_{0}=TS_{0}$ is the saddle-point and $\Omega_{2}=T\sum_{q}\ln\det[T\boldsymbol{\Gamma}^{-1}(q)]$ is the quadratic contribution. At $T_{c}$, while the former leads to the Fermi-Dirac (FD) distribution of an unbound (free) Fermi gas of $\uparrow$ and $\downarrow$ fermions, the latter contribution is typically split into $\mathcal{N}_{2}=\mathcal{N}_{bs}+\mathcal{N}_{sc}$  [22, 26]. While the bound-state contribution $\mathcal{N}_{bs}$ arises from the isolated poles of $\boldsymbol{\Gamma}(q)$, the continuum of two-particle excitations leads to the scattering contribution $\mathcal{N}_{sc}$ arising from the branch cut of the logarithm. Thus, $\mathcal{N}_{bs}=-T\sum_{q}\frac{\partial[\det\boldsymbol{\Gamma}^{-1}(q)]/\partial\mu}{\det\boldsymbol{\Gamma}^{-1}(q)}$ can be simplified considerably by noting that $\det\boldsymbol{\Gamma}^{-1}(q)\propto\Pi_{s}(\mathrm{i}\nu_{\ell}-\omega_{s\mathbf{q}}),$ where $\omega_{s\mathbf{q}}$ are the poles of $\boldsymbol{\Gamma}(q)$ determined by $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega_{s\mathbf{q}})=0.$ For instance, since $\partial\omega_{s\mathbf{q}}/\partial\mu=-2$ in the BEC limit, one can approximate $\mathcal{N}_{bs}=-T\sum_{\ell s\mathbf{q}}2/(\mathrm{i}\nu_{\ell}-\omega_{s\mathbf{q}})=2\sum_{s\mathbf{q}}f_{\mathrm{BE}}(\omega_{s\mathbf{q}}),$ after summing over the Matsubara frequencies, where $f_{\mathrm{BE}}(x)=1/(e^{x/T}-1)=[\coth\big{(}\frac{x}{2T}\big{)}-1]/2$ is the Bose-Einstein (BE) distribution of a bound Bose gas of Cooper pairs. On the other hand, the branch-cut contribution is given by $\mathcal{N}_{sc}=\sum_{\mathbf{q}}\int_{-\infty}^{+\infty}\frac{d\omega}{\pi}f_{\mathrm{BE}}(\omega)\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu},$ where $\delta(\mathbf{q},\omega)=-\mathrm{Arg}[\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega+\mathrm{i}0^{+})]$ is the argument of the the propagator [22, 26]. For instance, if we decompose $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega+\mathrm{i}0^{+})=R+\mathrm{i}I$ into its real and imaginary parts then $\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu}=(\dot{R}^{\mu}I-R\dot{I}^{\mu})/(R^{2}+I^{2}),$ where $\dot{R}^{\mu}=\partial R/\partial\mu$ and $\dot{I}^{\mu}=\partial I/\partial\mu.$ Since accurately computing $\mathcal{N}_{sc}$ is quite challenging, even in the context of much simpler continuum problems [22, 26], we next focus only on a dilute flat-band superconductor where $\mathcal{N}_{sc}$ is negligible for all $U\neq 0$, and can be omitted.
Thus, having a dilute flat-band superconductor in the pyrochlore lattice in mind, our self-consistency relations for $\mu$ and $T_{c}$ can be summarized as
11 11 $=\frac{U}{N}\sum_{n\mathbf{k}}\frac{\mathcal{X}_{n\mathbf{k}}}{2\xi_{n\mathbf{k}}},$ (11)
$F$ $=\frac{2}{N}\sum_{n\mathbf{k}}f_{\mathrm{FD}}(\xi_{n\mathbf{k}})+\frac{2}{N}\sum_{\mathbf{q}}f_{\mathrm{BE}}(\omega_{b\mathbf{q}}),$ (12)
where $0\leq F=\mathcal{N}/N\leq 2$ is the average filling of particles per lattice site and $f_{\mathrm{FD}}(x)=1/(e^{x/T}+1)=[1-\tanh\big{(}\frac{x}{2T}\big{)}]/2$ is the FD distribution. In Eq. (12), the prefactor of 2 in front of the FD distribution is due to spin degeneracy, while that of the BE distribution can be traced back to the presence of 2 particles in a two-body bound state. Since the flat bands appear at the bottom of the Bloch spectrum in the pyrochlore lattice, we emphasize that Eq. (12) is very accurate in the $F\ll 1$ limit for all $U\neq 0$, but it may still give qualitatively correct results up to the half-filling for the flat bands, i.e., up to $F\lesssim 0.5$. In addition, it is, by construction, accurate for other multiband lattices up to their half fillings, i.e., up to $F\lesssim 1$, in the $U/t\gg 1$ limit. In Eq. (12), it is sufficient to keep only one of the pole contributions, i.e., the one with the lowest energy, because poles with higher energies are expected to give negligible contributions due to the BE distribution. Note that, in the presence of time-reversal symmetry and uniform pairing, this is also equivalent to keeping the isolated pole of Eq. (3). Furthermore, in our numerical calculations, we implement the simplest approach and extract $\omega_{b\mathbf{q}}$ from the dispersion $E_{1\mathbf{q}}$ of the lowest-lying two-body bound-state branch as follows.
In general, all of the two-body bound states can be determined from the isolated poles of Eq. (2) exactly, i.e., by computing $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},E_{s\mathbf{q}})=0,$ after setting the thermal factors $\mathcal{X}_{n\mathbf{k}\sigma}\to 1$ and $\mu\to 0$ [5]. However, in the presence of time-reversal symmetry and uniform pairing, the lowest-lying branch can simply be determined from the isolated pole of Eq. (3), i.e., by computing $\Gamma_{0}^{-1}(\mathbf{q},E_{1\mathbf{q}})=0,$ after again setting the thermal factors $\mathcal{X}_{n\mathbf{k}}\to 1$ and $\mu\to 0$. It can be shown that the overall structure of the two-body bound-state branches $E_{s\mathbf{q}}$ resemble the underlying Bloch bands $\varepsilon_{n\mathbf{k}}$ but with the opposite sign of energy [6]. For instance, $E_{1\mathbf{q}}$ resembles to $-\varepsilon_{1\mathbf{k}}$ but with an effective pair hopping parameter $t_{b}$ and some energy offset depending on $U$. To determine these effective parameters, we first expand $\varepsilon_{1\mathbf{k}}=-6\bar{t}+a^{2}\bar{t}|\mathbf{k}|^{2}/8$ in the low $\mathbf{k}\to\mathbf{0}$ limit, and identify the relation between the hopping parameter and the effective mass $m_{1}$ of the unpaired spin-$\sigma$ particles as $\bar{t}=4/(a^{2}m_{1})$. This suggests that $t_{b}=4/(a^{2}M_{b})$ is the relation between the effective hopping parameter and effective mass $M_{b}$ of the lowest-lying two-body bound states.
In the low-$\mathbf{q}$ limit, we also recall that $E_{1\mathbf{q}}=E_{b}+\frac{1}{2}\sum_{ij}(M_{b}^{-1})_{ij}q_{i}q_{j}$  [15], where the energy offset $E_{b}$ is determined by $1=\frac{U}{N}\sum_{n\mathbf{k}}1/(2\varepsilon_{n\mathbf{k}}-E_{b}).$ Furthermore, very similar to that of the many-body bound states discussed in Sec. II.3, the inverse effective-mass tensor $(M_{b}^{-1})_{ij}=(M_{b}^{-1})^{\mathrm{intra}}_{ij}+(M_{b}^{-1})^{\mathrm{inter}}_{ij}$ of the lowest-lying two-body bound states is composed of an analogous conventional contribution $(M_{b}^{-1})^{\mathrm{intra}}_{ij}=\frac{1}{2D}\sum_{n\mathbf{k}}\ddot{\varepsilon}_{n\mathbf{k}}^{ij}/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}$ and an analogous geometric one $(M_{b}^{-1})^{\mathrm{inter}}_{ij}=\frac{1}{D}\sum_{n\mathbf{k}}g_{ij}^{n\mathbf{k}}/(2\varepsilon_{n\mathbf{k}}-E_{b})-\frac{1}{D}\sum_{n,m\neq n,\mathbf{k}}g_{ij}^{nm\mathbf{k}}/(\varepsilon_{n\mathbf{k}}+\varepsilon_{m\mathbf{k}}-E_{b}).$ It is important to emphasize that these two-body expressions are exact for any $U\neq 0$ in the pyrochlore lattice [6], and $(M_{b}^{-1})_{ij}=\delta_{ij}/M_{b}$ is diagonal as a consequence of uniform pairing. Thus, $E_{b}$ and $M_{b}$ can be determined self-consistently through these analytical expressions as functions of $U$, where $D=\sum_{n\mathbf{k}}1/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}.$ Once $E_{b}$ and $M_{b}$ are determined, we identify that
$\omega_{b\mathbf{q}}=E_{1\mathbf{q}}-2\mu=\varepsilon_{b\mathbf{q}}-\mu_{b},$ (13)
[…]